OpenStudy (anonymous):

For 2x - 3y = -10, when x = 4, What is the value of y?

OpenStudy (anonymous):

2(4)-3y=-10 8-3y=-10 -3y=-10-8 -3y=-18 divide by -3 on each side y=6 Do you get this?

OpenStudy (anonymous):

oh yeses so you just plug 4 in for x yes i understand thank you so much!! @cupcakelover2222

OpenStudy (anonymous):

Yup! sorry, just should've said that in the first place! Lol (: any more help for today, just tag me if im online!

OpenStudy (anonymous):

i will aha thanks (: @cupcakelover2222